OpenStudy (anonymous):

Calulate the slope of the line between points (7,-6) and (-3,-2) A. -2/5 B. -5/2 C. -1/13 D. -13

OpenStudy (katrinakaif):

You try; Use the equation and we can guide you through

Directrix (directrix):

A. -2/5 -4/10 = -2/5

OpenStudy (anonymous):

I thought so but wasnt sure

sam (.sam.):

\[Slope\frac{y _{2}-y _{1}}{x _{2}-x _{1}}\]

OpenStudy (katrinakaif):

Next time, post your answer first, and then we can correct you

OpenStudy (anonymous):

use the equation. . substitute the given points. .

OpenStudy (anonymous):

hey ma'am/sir,. . i know there is a moderator out there. . how can i change my profile name?

OpenStudy (anonymous):

I have no idea

Directrix (directrix):

@Vince ==> Let's move on.

OpenStudy (katrinakaif):

Currently there is no feature that enables you to change your username. (In fact, it would cause a confusion on Openstudy) The only other way until the feature becomes available is to create a new account; Take into consideration that you will loose all your medals and achievements as well.

sam (.sam.):

You need to learn otherwise you will fail the course

sam (.sam.):

There is one great site you can go http://www.khanacademy.org/ If you haven't heard of it.

sam (.sam.):

They have lots of math exercises and videos to guide you.

OpenStudy (anonymous):

@sam. is it the same with open study?

sam (.sam.):

No, open study can ask questions, but in http://www.khanacademy.org/ quite difficult to ask

OpenStudy (katrinakaif):

Openstudy is more of a generator for live guidance, similar to that of Yahoo Answers, Wiki, and even Tutorvista

sam (.sam.):

We're just helping him to learn the techniques, not to copy answers :)